# Towards Solving the Gilbert-Pollak Conjecture via Large Language Models

### v1

@@ B.5 Correctness Proof for Algorithm 2
    Otherwise, we have $\delta_{S^{-}}=1$. Since $N_{t^{*}}+\delta_{S^{+}}\leq 1$, inequality 5 becomes $\frac{\partial F}{\partial x}\leq N_{t*}+\delta_{S^{+}}-1\leq 0$
In all accepted cases, $\frac{\partial F}{\partial x}\leq 0$. Thus, Algorithm 2 provides a rigorous sufficient condition for monotonicity.
By combining the theoretical foundations from Section B.4 with the proof of correctness in this section, we establish the formal rigor of our entire verification pipeline. This ensures its reliability in handling both bounded and unbounded parameter regions, thus providing comprehensive support for the application of Theorem 6.
Appendix C Proofs of LLM-Generated Lemmas
In this section, we present the formal proofs for the geometric lemmas generated by the LLM agent during the iterative process. As discussed in Section E.2, the LLM proposes these lemmas based on the representative regime where $D$ is a regular point and the parameter $f$ is fixed at its boundary value $f=d$.
@@ Appendix C Proofs of LLM-Generated Lemmas
To apply these lemmas to the global verification problem (spanning all four cases defined in Section E.2), we adhere to the following transplantation policies:
@@
Lemma 22 (Step 1, Side AA).
If $c<1$, then there exists a regular point $E$ such that $|AE|\leq\max\{|AY|,|AZ|\}$.
@@
If $f<d$, then there exists a regular point $F$ such that $|DF|\leq\max\{|DZ|,|DQ|\}$.
Note that these 2 lemmas are the same as Corollary 33 and Corollary 34, respectively, when $D$ is a regular point. We restated them here for a clearer step-by-step proof.
@@ 3. Intersection Angle:
  • If $NT$ does not intersect $AW$, then should $T$ also be a Steiner point, the extension of the subsequent clockwise edge from $T$ must intersect either $XY$ or $YZ$. By Lemma  16, it follows that a trapped regular point, let us call it $E$, exists within the polygon $WXYZR$.
@@
The preceding sections have established a complete theoretical framework for both verifying (via Theorem 6, Section B.4 and the previous subsection) and discovering (via Section B.2) new verification functions. This section presents the mathematical outcomes of applying this framework. The addition of these newly discovered and rigorously proven lemmas to our verification set is what ultimately enables the improvement of the lower bound on the Steiner ratio.
Appendix D Algorithms and Implementation Details
D.1 Detail Algorithm for Reward Model
@@
    return False    end if   Extract region $B=\prod_{i=1}^{n}[l_{i},r_{i}]$ with $B\cap\text{Domain}\neq\emptyset$ from $Q$. Define $I_{\infty}\leftarrow\{i\mid r_{i}=+\infty\}$ and $V\leftarrow\text{Vertices}(B)$. Set $verified\leftarrow\textbf{False}$.  for $F\in\mathcal{F}$ do     // Check monotonicity conditions if $F$ is monotonically decreasing in all $i\in I_{\infty}$ then        if $\max_{v\in V}F(v,\rho)\leq 0$ then           $verified\leftarrow\textbf{True}$          end if               end if         end for  if not verifiedverified then     $k^{*}\leftarrow\textstyle\arg\max_{i}(\text{if }i\in I_{\infty}\\ \text{ then }1/l_{i}\text{ else }r_{i}-l_{i})$$m\leftarrow\text{if }k^{*}\in I_{\infty}\text{ then }2\cdot l_{k^{*}}\text{ else }(l_{k^{*}}+r_{k^{*}})/2$  Split $B$ along dimension $k^{*}$ at $m$ into sub-regions $B_{left}$ and $B_{right}$  Push $B_{left}$ and $B_{right}$ to $Q$    end if   end while return True
@@
    Monotonicity Check for Unbounded Regions. (Line 7) To handle regions extending to $\infty$, we verify that the verification function $f$ is monotonically non-increasing along the unbounded dimension. This ensures that the maximum value is attained at the finite lower bound. The implementation details of the monotonicity check are provided in Appendix Section B.4.
@@
D.2 Monotonicity Verification Algorithm
@@
Output: True if $\frac{\partial F}{\partial x}\leq 0$ is guaranteed; False otherwise
@@
for each pair of adjacent terminals $(u,v)$ in $S^{+}$ do
@@
Based on the lemmas above, we define the procedure to verify $\frac{\partial F}{\partial x}\leq 0$, which is presented in Algorithm 2.
D.3 Strategies for Handling Complex Splits
D.3.1 Methods for Handling Splits with Implicit Regular Points
When a split involves implicit regular points and utilizes edges connected to them (referred to as implicit edges), the validity condition and length functions for these edges are complex. Consequently, we only perform monotonicity checks for variables with respect to which the implicit edge lengths are constant. For these variables, the above procedure remains valid. For other variables, we do not regard the splitting function as having monotonicity with respect to these variables.
Specifically, under the settings in Sections E and C, we employ the following strategy:
  • Splits involving the implicit points $E$: For variables $s$, $c$, and $e$, we do not assume monotonicity. For other variables, Algorithm 2 can be directly applied.
  • Splits involving the implicit points $F$: For variables $d$, $c$, $f$ (and $u$ when $D$ is Steiner), we do not assume monotonicity. For other variables, Algorithm 2 can be directly applied.
@@ D.3.1 Methods for Handling Splits with Implicit Regular Points
D.3.2 Methods for Handling Splits with 4-Point Steiner Tree
When a split comprises an $S^{+}$ that connects four terminals, the monotonicity of variable $f$ can be verified using Algorithm 2, as established in Appendix B.4.2. For all other variables, monotonicity is not assumed.
@@ D.3.2 Methods for Handling Splits with 4-Point Steiner Tree
D.4 Enumeration of Valid Splits
To generate a rich candidate set of splitting functions, we perform an algorithmic enumeration of all topologically valid splits. The procedure is formally defined in Algorithm 3.
@@ D.4 Enumeration of Valid Splits
Appendix E Experimental Setup and Additional Results
E.1 Baseline Verification Functions
Figure 11: Sample of Subtree: $D$ is regular
Du & Hwang (1983) Lemma 2
Lemma 32 (Du & Hwang (1983) Lemma 2).
Corollary 33.
Corollary 34.
If $f\leq d$, there must exist an implicit regular point $F$ such that its distance to the terminal is bounded. The bound depends on whether $D$ is a regular point or a Steiner point: If $D$ is a regular point, $|DF|\leq\max\{|DZ|,|DQ|\}$. If $D$ is a Steiner point, $|UF|\leq\max\{|UY|,|UZ|,|UQ|\}$.

### v2

@@ D.1 Baseline Verification Functions
    The set of admissible edges in $t^{*}$ is augmented to include the implicit connections $(A,E)$ and $(D,F)$ (or $(V,F)$). The length functions for these edges are conditional:
    • Edge $(A,E)$:
    • Edge $(D,F)$ (when $D$ is a regular point):
    • Edge $(V,F)$ (when $D$ is a Steiner point):
      $L(V,F)=\begin{cases}\max\{|VY|,|VZ|,|VQ|\}&f\leq d\\ +\infty&f>d\end{cases}$
Impact on Monotonicity. When a split utilizes implicit edges $(A,E)$, $(D,F)$, or $(V,F)$, their length bounds depend on local geometric parameters. Establishing the required non-increasing property for these variables is non-trivial. Consequently, our system performs the monotonicity checks (as detailed in Appendix F.2) only for variables with respect to which the implicit edge lengths are constant.
Explanation of Steiner Ratio greater than $0.824$ for Baseline Settings. Though we only cited technical lemmas from Du & Hwang (1983) as the baseline, and that work only claimed a proof for a Steiner ratio lower bound of $\rho=0.8$, our initial ratio reached $0.8282$, even surpassing the previous best-known lower bound of $0.824$ by Chung & Graham (1985). This improvement is due to (1) prior work considered smaller topological structures. Our baseline setting directly considered topologies with a maximum chain length of $4$, as shown in Figure 10, while previous human experts, constrained by computational capabilities, only considered topologies with a maximum chain length of $3$. This allowed us to more accurately cover the parameter space, resulting in a higher lower bound; (2) prior work relied on manual proofs considering a relatively limited set of verification functions, whereas we leveraged computational mathematics to exhaustively enumerate all possible splits; (3) previous work employed looser scaling techniques, while we utilized more precise computational algorithms, eliminating gaps caused by arithmetic tricks.
D.2 Problem Decomposition and Search Space Reduction
  1. 1.
    Case 1: $D$ is a regular point, and $f\in[0,d]$ ($F$ exists).
  2. 2.
    Case 2: $D$ is a regular point, and $f\in(d,+\infty)$ ($F$ does not exist).
  3. 3.
    Case 3: $D$ is a Steiner point, and $f\in[0,d]$ ($F$ exists).
  4. 4.
    Case 4: $D$ is a Steiner point, and $f\in(d,+\infty)$ ($F$ does not exist).
Reduction Strategy for Unbounded $f$. For Cases 2 and 4, where $f>d$, the implicit connection $(D,F)$ (or $(U,F)$ and $(V,F)$) is invalid and thus excluded from the connector graph $t^{*}$. In these scenarios, we apply a strict filtering policy: we only admit splits whose splitting functions are monotonically non-increasing with respect to $f$.
Hence, in the region $f\geq d$, it suffices to verify the boundary $f=d$. For each function $F$, define $F^{\partial}$ as the reduced function obtained from $F$ by setting $f=d$ and removing $f$ from the input variables. The verification target is thus reduced to
$\max_{{\bm{w}}\in\mathcal{W}^{\partial}}\min_{F\in\mathcal{F}}F^{\partial}({\bm{w}})\leq 0,$
where $\mathcal{W}^{\partial}$ denotes the reduced parameter space obtained by removing the $f$-coordinate from $\mathcal{W}$, so that $\mathcal{W}^{\partial}=[0,+\infty)^{n-1}$.
Note that $F^{\partial}$ preserves both convexity and coordinatewise non-increasingness of $F$. In particular, its non-increasingness with respect to $d$ follows because increasing $d$ simultaneously increases the original $d$-entry and the substituted $f$-entry, and $F$ is non-increasing in both variables.
  1. 1.
  2. 2.
Appendix E Proofs of LLM-Generated Lemmas
In this section, we present the formal proofs for the geometric lemmas generated by the LLM agent during the iterative process. As discussed in Section D.2, the LLM proposes these lemmas based on the representative regime where $D$ is a regular point and the parameter $f$ is fixed at its boundary value $f=d$.
@@ Appendix E Proofs of LLM-Generated Lemmas
To apply these lemmas to the global verification problem (spanning all four cases defined in Section D.2), we adhere to the following transplantation policies:
@@
Lemma 27 (Step 1, Side $A$).
If $c\leq 1$, then there exists a regular point $E$ such that $|AE|\leq\max\{|AY|,|AZ|\}$.
@@
If $f\leq d$, then there exists a regular point $F$ such that $|DF|\leq\max\{|DZ|,|DQ|\}$.
Note that these 2 lemmas are the same as Corollary 25 and Corollary 26, respectively, when $D$ is a regular point. In particular, they are not model-generated lemmas. We restated them here for a clearer step-by-step proof.
@@ 3. Intersection Angle:
  • If $NT$ does not intersect $AW$, then should $T$ also be a Steiner point, the extension of the subsequent clockwise edge from $T$ must intersect either $XY$ or $YZ$. By Lemma  18, it follows that a trapped regular point, let us call it $E$, exists within the polygon $WXYZR$.
@@
Appendix F Algorithms and Implementation Details
F.1 Detail Algorithm for Reward Model
@@
    return False    end if   Extract region $B=\prod_{i=1}^{n}[l_{i},r_{i}]$ with $B\cap\text{Domain}\neq\emptyset$ from $Q$. Define $I_{\infty}\leftarrow\{i\mid r_{i}=+\infty\}$ and $V\leftarrow\text{Vertices}(B)$. Set $verified\leftarrow\textbf{False}$.  for $F\in\mathcal{F}$ do     // Check monotonicity conditions if $F$ is monotonically decreasing in all $i\in I_{\infty}$ then        if $\max_{v\in V}F(v,\rho)\leq 0$ then           $verified\leftarrow\textbf{True}$          end if               end if         end for  if not verifiedverified then     $k^{*}\leftarrow\displaystyle\operatorname*{arg\,max}_{i}\left(\text{if }i\in I_{\infty}\text{ then }1/l_{i}\text{ else }r_{i}-l_{i}\right)$$m\leftarrow\text{if }k^{*}\in I_{\infty}\text{ then }2\cdot l_{k^{*}}\text{ else }(l_{k^{*}}+r_{k^{*}})/2$  Split $B$ along dimension $k^{*}$ at $m$ into sub-regions $B_{left}$ and $B_{right}$  Push $B_{left}$ and $B_{right}$ to $Q$    end if   end while return True
@@
    Monotonicity Check for Unbounded Regions. (Line 7) To handle regions extending to $\infty$, we verify that the verification function $f$ is monotonically non-increasing along the unbounded dimension. This ensures that the maximum value is attained at the finite lower bound. The implementation details of the monotonicity check are provided in Appendix F.2.
@@
F.2 Monotonicity Verification Algorithm
The vertex-checking strategy of Theorem 6 is not directly applicable to unbounded parameter regions. To address this case, we must first verify the function’s monotonicity along any unbounded dimensions, thereby reducing the problem to one on a bounded domain. Based on the lemmas in Appendix C.4 (Lemmas 21, 22, and 23), we define the procedure to verify $\frac{\partial F}{\partial x}\leq 0$, which is presented in Algorithm 2.
@@ F.2 Monotonicity Verification Algorithm
Output: True if $\frac{\partial F_{\tau}}{\partial x}\leq 0$ is guaranteed; False otherwise
@@
for each pair of distinct terminals $(u,v)$ in $S^{+}$ do
@@
Here, $\delta_{S^{-}}$ indicates whether $e$ belongs to $S^{-}$, $N_{t^{*}}$ is the upper bound in Lemma 21, and $\delta_{S^{+}}$ indicates whether $e$ lies on the path between some pair of distinct terminals in $S^{+}$. These quantities are then combined to certify whether $\frac{\partial F}{\partial x}\leq 0$ is guaranteed.
Proposition 37 (Correctness of Algorithm 2).
Suppose that either
  1. 1.
    $S^{+}$ connects at most three terminals, or
  2. 2.
    the parameter under consideration is $x=f$.
If Algorithm 2 returns True, then
$\frac{\partial F_{\tau}}{\partial x}\leq 0.$
Proof.
$\frac{\partial F_{\tau}}{\partial x}=\rho\frac{\partial L_{t^{*}}}{\partial x}+\frac{\partial L_{S^{+}}}{\partial x}-\frac{\partial L_{S^{-}}}{\partial x}.$
We analyze the bounds using the variables computed in Algorithm 2:
  • By Lemma 21, $\frac{\partial L_{t^{*}}}{\partial x}\leq N_{t^{*}}$.
  • For the $S^{+}$ term, if $\delta_{S^{+}}=0$, then $e$ lies on no path between any pair of terminals in $S^{+}$, so $L_{S^{+}}$ is independent of $x$ and $\frac{\partial L_{S^{+}}}{\partial x}=0$. If $\delta_{S^{+}}=1$, then $\frac{\partial L_{S^{+}}}{\partial x}\leq 1$ follows from Lemma 22 when $S^{+}$ connects at most three terminals, and from Lemma 23 when $x=f$. Therefore, in either case, $\frac{\partial L_{S^{+}}}{\partial x}\leq\delta_{S^{+}}$.
Substituting these into the derivative:
$\frac{\partial F_{\tau}}{\partial x}$ $\leq\rho\cdot N_{t^{*}}+\delta_{S^{+}}-\delta_{S^{-}}\leq N_{t^{*}}+\delta_{S^{+}}-\delta_{S^{-}}.$ (5)
Whenever Algorithm 2 returns True, either $N_{t^{*}}+\delta_{S^{+}}\leq 1$ and $\delta_{S^{-}}=1$, or $\delta_{S^{-}}=N_{t^{*}}=\delta_{S^{+}}=0$. In either case, equation 5 implies
$\frac{\partial F_{\tau}}{\partial x}\leq 0.$
Thus, Algorithm 2 provides a rigorous sufficient condition for monotonicity. ∎
F.3 Strategies for Handling Complex Splits
F.3.1 Methods for Handling Splits with Implicit Regular Points
When a split involves implicit regular points and utilizes edges connected to them (referred to as implicit edges), the splitting function cannot be evaluated directly, since its value depends on the locations of these implicit regular points. We therefore analyze instead the monotonicity of the verification functions generated by the Trapped Regular Point Lemma. We only perform monotonicity checks for the baseline verification functions; verification functions generated by other lemmas are not regarded as monotone.
For a baseline verification function, Algorithm 2 is applied only to variables with respect to which the implicit edge lengths are constant. For these variables, Proposition 37 remains valid. If an implicit edge length depends on a variable, we do not regard the corresponding verification function as monotone with respect to that variable. Under the settings in Section D, the variables with respect to which each implicit edge length is not constant are listed in Table 2.
Implicit edge Nonconstant variables
$(A,E)$ $c,s$
$(V,F)$ $v,d,c,f$
$(U,E)$ $c,d,s,u,e$
$(V,E)$ $c,d,s,v,e$
When $D$ is a Steiner point
Implicit edge Nonconstant variables
$(A,E)$ $c,s$
$(D,F)$ $d,c,f$
$(D,E)$ $c,d,s,e$
When $D$ is a regular point
Table 2: Variables with respect to which each implicit edge length is not constant.
@@ F.3.1 Methods for Handling Splits with Implicit Regular Points
F.3.2 Methods for Handling Splits with 4-Point Steiner Tree
When a split comprises an $S^{+}$ that connects four terminals, the monotonicity of variable $f$ can be verified using Algorithm 2, as established in Proposition 37. For all other variables, monotonicity is not assumed.
@@ F.3.2 Methods for Handling Splits with 4-Point Steiner Tree
F.4 Enumeration of Valid Splits
To generate a rich candidate set of splitting functions, we perform an algorithmic enumeration of topologically valid splits. The procedure is formally defined in Algorithm 3.
@@ F.4 Enumeration of Valid Splits
The algorithm enumerates splittings in three phases.
First, it chooses every non-empty subset $V^{*}\subseteq V_{exp}$ and enumerates edge sets $S^{-}$ such $S^{-}$ contains all edges incident to vertices in $V^{*}$.
Second, let
$G^{\prime}=(V(G),E(G)\setminus S^{-}).$
The algorithm identifies all connected components of $G^{\prime}$ that contain at least one vertex from $W\setminus V^{*}$, where $W$ consists of all explicit regular points together with the boundary points. Since every path from a point in $V\setminus V_{exp}$ to the local graph $G$ enters through a boundary point, the components containing $W\setminus V^{*}$ capture all locally relevant residual pieces. The algorithm then chooses one representative vertex from each such component and, for each resulting tuple of representatives, takes $S^{+}$ to be the SMT on those vertices.
Finally, the algorithm enumerates candidate forests $t^{*}$ for reconnecting $V^{*}$ to the remaining regular points. To test whether a candidate has the required topology, all vertices in $V_{conn}\setminus V^{*}$ are contracted into a single super-node $\Omega$. The candidate is accepted if the resulting multigraph is a spanning tree on $V^{*}\cup\{\Omega\}$, with self-loops and parallel edges taken into account. This guarantees that every removed vertex is connected to the outside, while no two outside vertices are connected through $t^{*}$.
Thus, the algorithm systematically generates a broad set of valid splittings by exploring different choices of $V^{*}$, $^{-}, S+S^{+}$, and $t^{*}$.
Appendix G LLM Agent and Prompts
G.1 Prompt for Searching Trapped Regular Point Lemmas
[…]
G.2 Prompt for Searching Valid 4-Point Steiner Tree Lemmas
Prompt for Searching Valid 4-Point Steiner Tree Lemmas PROMPT START Role: You are an expert Computational Geometer and Mathematician. Objective: Derive a simplified, robust condition $P(b,c,d,s,e)$ that guarantees the existence of a specific Steiner Tree. Constraint: The condition you find must be “Axis-Parallel Segment Closed” (APSC). Tools: You have access to a Mathematica execution environment and a Python execution environment. PHASE 1: Geometric Definitions Construct the following points in the 2D plane based on these vector instructions. A: $(0,0)$ X: From A, move length $1$ at angle $0^{\circ}$. Y: From X, move length $s$ at angle $60^{\circ}$. B: From X, move length $b$ at angle $60^{\circ}$. D: From Y, move length $d$ at angle $0^{\circ}$. P: From Y, move length $c$ at angle $120^{\circ}$. Q: From P, move length $d$ at angle $60^{\circ}$. R: From P, move length $e$ at angle $180^{\circ}$. The domain of the variables is $b\geq 0\;\&\&\;c\geq 0\;\&\&\;d\geq 0\;\&\&\;s\geq 0\;\&\&\;e\geq 0$. Note: Although Points X, Y, P are used for construction, your final task focuses on the Steiner Tree for points {A}, {B}, {C}, {D}. PHASE 2: The Existence Logic (Type {A}{B}-{C}{D}) You are investigating the existence of an ({A}{B})-({C}{D}) type Steiner Tree. For this specific type, define the auxiliary points $U$ and $V$ as follows: 1. Point U: Rotate point {B} $60^{\circ}$ counter-clockwise around {A}. 2. Point V: Rotate point {D} $60^{\circ}$ counter-clockwise around {C}. The Exact Existence Conditions ($E_{exact}$) are: 1. Convexity: The quadrilateral formed by {A, B, C, D} must be convex. 2. Internal Angles: The angles $\angle UCD$, $\angle UDC$, $\angle VAB$, and $\angle VBA$ must all be $\leq 120^{\circ}$. 3. Crossing Angle: Let $\theta$ be the angle between diagonal vectors $\vec{AC}$ and $\vec{BD}$. The condition is $\theta\leq 120^{\circ}$. Relaxation Hint: You may relax this third condition to $\theta\leq 90^{\circ}$ (which is $\vec{AC}\cdot\vec{BD}\geq 0$) if it helps simplify the math. The condition for “convex quadrilateral” can be formulated as follows: For any three adjacent points x, y, z (in clockwise order), the cross product of vectors xy and yz is less than or equal to zero. The condition that the angle between vectors u and v is less than or equal to $120$ degrees can be expressed using the cosine rule or the dot product of vectors, namely: vectors u and v form an angle $\leq 120^{\circ}$ iff $2\,\mathrm{dot}(u,v)+(|u|\,|v|)\geq 0.$ You may relax it to $\mathrm{dot}(u,v)\geq 0$ if possible. PHASE 3: Deriving the Relaxed Condition P The exact conditions above are too complex. You must find a sufficient condition $P(b,c,d,s,e)$ such that $P\implies E_{exact}$. Requirements for Condition P (APSC Property): Definition: A condition $P$ satisfies APSC if, for any line parallel to one of the coordinate axes, the set of points within the domain (variables $\geq 0$) that satisfy the condition forms a single connected interval. To satisfy APSC, $P$ must be the conjunction of inequalities from the following allowed categories only: 1. Linear or Multilinear: Inequalities that are linear in each variable separately. 2. Quadratic: Inequalities in the form $Ax^{2}+Bx+C\geq 0$ (or $\leq 0$), subject to monotonicity rules explained in the prompt. 3. Functional Boundaries: Inequalities in the form $e>f(b,c,s,d)$ or $e<f(b,c,s,d)$, where monotonicity of $f$ must be verified. Coverage Requirement: Your condition $P$ must be general (not hardcoded numbers), but it must strictly cover the following “Test Box”: {Bottleneck} Important: You MUST verify the APSC Property in the ENTIRE domain, not in the Test Box. PHASE 4: Verification Protocol You must verify your work using Mathematica logic. Step A: Write Mathematica code to compute the exact expressions for the 6 existence inequalities (Convexity + Angles). Step B: Define your relaxed condition condP. Step C: Verify Implication. You must run a command equivalent to this logic (check each inequality separately): Reduce[ !<ExactCondition_i> && condP && b>=0 && c>=0 && d>=0 && s>=0 && e>=0, {vars} ] If this returns False, your condition is valid. If it returns True or an expression, your condition is invalid. Tips: Reformulate $2\,\mathrm{dot}(u,v)+(|u|\,|v|)\geq 0$ into $\mathrm{dot}(u,v)\geq 0\;||\;4\,\mathrm{dot}(u,v)^{2}\leq|u|^{2}|v|^{2}$ for faster speed when verifying. Constraint: Do not put all checks in one Reduce call. It will time out. Check each inequality individually against $P$. Important: You MUST verify the implication in the ENTIRE domain, not in the Test Box. PHASE 5: Output Requirements 1. Detailed Derivation: Show how you calculated coordinates and how you derived the relaxed inequalities. 2. APSC Proof: Provide a logical argument explaining why your condition $P$ satisfies the APSC property, specifically addressing the Linearity/Monotonicity requirements for each part of your condition. 3. Mathematica Output: Paste the code and the output confirming the False result for the implication checks. 4. C++ Implementation: Provide a C++ code block containing these exact functions: // Returns true if your condition \$P\$ is met. bool steiner_cond(double b, double c, double d, double s, double e) { // Your condition here } // Returns the total length of the ({A}{B}-{C}{D}) type Steiner tree, which is exactly |UV|, the distance between U and V. double steiner_length(double b, double c, double d, double s, double e) { // Return the distance between U and V } Important: Do NOT use the Test Box values inside the C++ code. The code must be a general algebraic formula. If you fail to find a valid condition after rigorous attempts, return false and $0.$ — END OF PROMPT —
Appendix H A Simplified Version for Mathematicians
If you are interested solely in the mathematical results of this work and are not concerned with the content related to LLMs, you may proceed directly to this section. We will present the complete logical structure of the mathematical proofs in the most concise manner possible.
H.1 Preliminaries
H.1.1 Definitions and Notations
We first introduce the necessary definitions and notations.
Definition 1 (Minimum Spanning Tree). Consider a set $V$ of $n$ points in the Euclidean plane $\mathbb{R}^{2}$. A spanning tree on $V$ is a connected, acyclic graph with vertex set $V$. When the length of each edge is defined as the Euclidean distance between its endpoints, a spanning tree that minimizes the total length is called a Minimum Spanning Tree.
Definition 2 (Steiner Minimal Tree). Consider a set $V$ of $n$ points in the Euclidean plane. The shortest network interconnecting all points in $V$, where the length of each edge is measured by Euclidean distance, is necessarily a tree, referred to as a Steiner Minimal Tree. A Steiner Minimal Tree may contain auxiliary vertices not in $V$. These additional vertices are called Steiner points, while the points in $V$ are referred to as regular points.
Then we formally define the Steiner ratio and the Steiner ratio conjecture.
Definition 3 (Steiner Ratio). The Steiner ratio is defined as the infimum of the ratio of the length of the Steiner Minimal Tree to the length of the Euclidean Minimum Spanning Tree over all finite sets of points $V$:
$\rho_{\text{Steiner}}=\inf_{V}L_{S}(V)/L_{m}(V),$
where $L_{S}(V)$ and $L_{m}(V)$ denote the lengths of Steiner Minimal Tree and Minimum Spanning Tree, respectively.
H.1.2 Basic Properties
Let $V$ be a finite set of points in the Euclidean plane. We denote the lengths of the Steiner Minimal Tree (SMT) and the Minimum Spanning Tree (MST) on $V$ as $L_{S}(V)$ and $L_{m}(V)$, respectively. For an arbitrary geometric graph $G$, let $L_{G}$ denote the sum of the Euclidean lengths of its edges.
We first recall some fundamental geometric properties of SMTs.
Lemma 11 (Gilbert & Pollak (1968), Basic Properties of SMTs). Let $S$ be a Steiner Minimal Tree for a set $V$. The following properties hold: